a rectangle table is 4 times long as it is wide. if the area is 100ft, find the length and width of the table
Answer by kirti251(6)

Here rectangle table is 4 time long as its width & area is 100ft so
Guage width is x so lengh is 4x so
x * 4x=100
4x*x=100
x*x=25 so x=5